A loan of P5,000 is made for a period of 15 months, at a simple interest rate of 15%, what future amount is due at the end of the loan period?



1
Expert's answer
2021-09-23T08:54:41-0400

"Soln,"


Loan amount"=P5,000"

Period"= 15" months "=1.25 years"

Interest rate"=15 \%" "=0.15"


"A=P(1+rt)""=P(1+rate\times\ time)"

"A=P5000(1+0.15(15\/12)months)"

"A=P5000(1+0.1875)"

"=P5000(1.1875)"

Amount due "=5937.50"